Accueil > Sites heberges > CFHTLS-Deep > Recent updates

 

Recent updates

Par Genevieve Soucail - 23/10/2008

 

Recent updates

(last modification : October 23, 2008)


Observations are closed for the CFHTLS-Deep survey 


Summary of last runs (2007-2008 only)

  • Observations of D2 and D3 are closed and are finished for the CFHTLS-Deep. See the D2 and D3 web pages for details
  • Summary of queue run 08AQ09 : 27 May - 12 June
    • A very good run, with only 1.5 night lost to weather and 2 nights with poor seeing. The rest of the time had good or excellent observing conditions.
    • D2 observed for 2.2 hours in g’, r’, i’ and z’.
    • D3 observed for 8.6 hours in 5 filters.
  • Summary of queue run 08AQ07 : 30 Apr. - 14 May
    • A very good run, with only 1.5 night lost to weather and 2 nights with difficult conditions. The rest of the time had good or excellent observing conditions.
    • D2 observed for 4.4 hours in g’, r’, i’ and z’.
    • D3 observed for 8.8 hours in g’, r’, i’ and z’.
  • Summary of queue run 08AQ05 : 28 Mar. - 14 Apr.
    • A mediocre run, with 3 full nights lost to weather and 4 additional half-nights. The rest of the time had to face with clouds, humidity, and poor to variable seeing along the nights, making the observations difficult and not optimal. Finally a rather good amount of LS data taken yet.
    • D2 observed for 7.9 hours in g’, r’, i’ and z’.
    • D3 observed for 8.8 hours in 5 filters.
  • Summary of queue run 08AQ03 : 27 Feb. - 14 Mar.
    • A rather good run, although some time was lost at the beginning of the run due to thick cirrus ( 1/2 night), and on average the seeing was higly variable and sometimes too high for observing.
    • D2 observed for 8.1 hours in g’, r’, i’ and z’.
    • D3 observed for 9.5 hours in 5 filters.
  • Observations of D1 are closed and are finished for the CFHTLS-Deep. See the D1 web page for details
  • Summary of queue run 08AQ01 : 31 Jan. - 16 Feb.
    • A difficult start for the semester. 11.5 nights lost to weather over the run. No technical problems.
    • D1 observed for 0.7 hours only in i’.
    • D2 observed for 4.8 hours in g’, r’, i’ and z’.
    • D3 observed for 5.4 hours in 5 filters.
  • Summary of queue run 07BQ11 : 02 Jan. - 16 Jan.
    • A very good run compared to the previous one ! No more than 2 nights lost to strong absorption distributed over the run.
    • D1 observed for 8.7 hours in 5 filters.
    • D2 observed for 9.2 hours in g’, r’, i’ and z’.
    • D3 started for 1.2 hours in r’ and i’.
  • Summary of queue run 07BQ10 : 03 Dec. - 17 Dec.
    • A very poor run with 12 nights out of 14 lost to weather : snow, fog, bad seeing and strong absortion
    • D4 observed for 1.2 hours in r’ and i’.
    • D1 observed for 2.1 hours in g’, r’, i’ and z’.
  • Observations of D4 are closed and are finished for the CFHTLS-Deep. See the D4 web page for details
  • Summary of queue run 07BQ08 : 02 Nov. - 15 Nov.
    • A rather good run, with 4.5 nights list to weather in total. No technical problems.
    • D4 observed for 8 hours in g’, r’, i’ and z’.
    • D1 observed for 12.4 hours in 5 filters.
    • D2 started for 2.2 hours in g’, r’ and i’.
    • The new i’ filter is on the sky. But these new data will not be included in the final stacks of the Deep.
  • Summary of queue run 07BQ06 : 04 Oct. - 17 Oct.
    • A good run finally, after a dififcult start (2 nights lost to weather), and about 2 equivalent nights plagued with cirrus and/or poor seeing.
    • D4 observed for 9.5 hours in u*, g’, r’ and z’.
    • D1 observed for 7.3 hours in g’, r’ and z’.
  • Summary of queue run 07BQ06 : 04 Oct. - 17 Oct.
    • A good run finally, after a dififcult start (2 nights lost to weather), and about 2 equivalent nights plagued with cirrus and/or poor seeing.
    • D4 observed for 9.5 hours in u*, g’, r’ and z’.
    • D1 observed for 7.3 hours in g’, r’ and z’.
    • i’ filter still missing, but was back on the sky for the last night of th run (focus tests)
  • Summary of queue run 07BQ04 : 04 Sep. - 17 Sep.
    • An average run, with 4 nights lost to humidity, clouds and fog, plus a few technical problems. Decent seeing for most of the rest of the run, except a few hours.
    • D4 observed for 6.9 hours in g’, r’ and z’.
    • D1 observed for 7.3 hours in u*, g’, r’ and z’.
    • i’ filter still missing
  • Summary of queue run 07BQ02 : 06 Aug. - 20 Aug.
    • A very good first half of the run until Aug. 13 when hurricane Flossie approached Hawaii. Fortunately it rapidly retrograded and hit the island with less damage than expected. 3 nights lost to Flossie, followed by the rest of the run with only a few windows of observations with decent seeing.
    • D4 observed for 7.2 hours in u*, g’, r’ and z’.
    • D1 observed for 6.3 hours in u*, g’, r’ and z’.
    • i’ filter still missing
  • Summary of queue run 07AQ12 : 20 Jul. - 30 Jul.
    • MegaCam back on the telescope on July 17 for several nights of engeneering and extensive tests (a bit longer than expected partly because of foggy nights). Start of science observations on July 23, until July 30 with a full Moon at the end. A few hours with mediocre seeing.
    • D4 observed for 5.2 hours in g’, r’ and z’.
  • Observations of D3 are closed (Jul. 2007). See the D3 web page for details
  • Jun 17 : serious MegaCam failure with the loss of the i’ filter. All observations with MegaCam stopped. A new i’ filter has been ordered and should be available hopefully for the October run.
  • Summary of queue run 07AQ09 : 30 May - 18 Jun.
    • MegaCam back on the telescope after repair on May 30, even with a Moon at maximum ! A very good and long run, with only one night lost to clouds. Until June 17 when MegaCam crashed again and the run was stopped.
    • D3 observed for 12.9 hours (with 2.7 hours spent in u*)
    • D4 observed for 1.3 hours
  • Observations of D2 are closed (May 18, 2007). See the D2 web page for details
  • Summary of queue run 07AQ07 : 08 May - 17 May
    • Not a very good run : about 4.5 nights lost to thick cirrus and high humidity. Then during the night of May 16, a serious problem occurred with MegaCam when the jukebox crashed into the shutter. The camera had to go off the telescope for repair. This ended the run prematurely.
    • D2 observed for 5.5 hours
    • D3 observed for 3.8 hours in 4 filters
  • Summary of queue run 07AQ05 : 06 Arp. - 23 Apr.
    • A very good run, with only the last night lost to ice. About 4 nights with limited observations due to bad seeing ou cirrus. Again a good amount of CFHTLS data. No technical problems.
    • D2 observed for 10.4 hours
    • D3 observed for 11.9 hours in 5 filters (with 2.7 hours spent in u*)
  • Summary of queue run 07AQ03 : 11 Mar. - 27 Mar.
    • A difficult start for the run (5 nights lost to high humidity). Then the rest of the run was good to very good. Good amount of CFHTLS data. No technical problems.
    • D2 observed for 6.8 hours
    • D3 observed for 9.2 hours in 5 filters
  • Observations of D1 are closed (Feb. 22, 2007). See the D1 web page for details
  • Summary of queue run 07AQ02 : 08 Feb. - 22 Feb.
    • A rather good run for winter time. 3 nights lost to weather and 4 additional ones with poor observing conditions (bad seeing or clouds). No technical problems.
    • D1 observed for 1 hour in r’ and i’
    • D2 observed for 4.4 hours
    • D3 observed for 6.3 hours in 5 filters
  • Summary of queue run 06BQ10 : 11 Jan. - 25 Jan.
    • A very good run indeed ! except the first night lost to technical problems during the focus exchange. Some nights with variable seeing.
    • D1 observed for 6.6 hours with 1h in u*
    • D2 observed for 3.8 hours
    • D3 started with 7.7 hours

previous runs

 

 

[Page Précédente] [Dans la même rubrique] [Sommaire]

 

 

Site du CNRS Logo OMP
Logo UPS
Logo INSU-CNRS

Annuaire

Rechercher

Sur ce site



Bibliotheque

Laboratoire Astrophysique de Toulouse - Tarbes (UMR5572)

CNRS (Midi-Pyrénées)

Univ. Paul Sabatier

Liens utiles

Accueil Imprimer Contact mail Plan du site Crédits